InTheSea:

A 10.0 Ω resistor is in parallel with a 20.0 Ω resistor and supplied by a 9.00 V battery. What is the total current passing through the circuit? 1.35 A 0.38 A 1.13 A 6.00 A

InTheSea:

I think the answer is 1.35A, but I'm unsure of how that's the answer @Shadow

Shadow:

\[I_{total} = \frac{ V_{total} }{ R_{total}}\] \[R_{total} = (R^{-1} + R^{-1})^{-1}\] \[R_{total} = (10^{-1} + 20^{-1})^{-1} = \frac{ 20 }{ 3 } = 6.7 \Omega\] \[I_{total} = \frac{ 9 }{ \frac{ 20 }{ 3 } } = 1.35 A\] You were correct :)

Shadow:

@InTheSea